Currently browsing tag



TEM is back online

July 9, 2013. I “found” the electron beam and dialed in a good alignment on the TEM today and it is working fine. You can use it now. Note that the lab temp and humidity are such that we will begin using the liquid nitrogen cold traps. I will show …


TEM status – online

April 10, 2013. After the repair, I have tested the TEM and found beam current present each time.  I’m cautiously optimistic that it is repaired.  Please use the scope and let me know right away if you see problems.


TEM status

October 2, 2012. The TEM problem seems to be in remission so, I am releasing the TEM for routine use. Note that the dark current values have changed since we opened the HV tank. This will affect the maximum beam current values. IMPORTANT NOTE – The maximum kV is 200! …


TEM update

August 21, 2012. The TEM electron beam problem appears to be repaired. I have been able to achieve a stable 200kV beam ~10 times now so I believe the problem is repaired. I will work to achieve 300kV over the next few weeks. There is one problem that remains to …